simplify write final answer in the lowest terms
18/35+22/25

Answer by lenny460(1073)

18/35 + 22/35
35 * 25 = 875
875/35 x 18 = 450
875/25 x 22 = 770
450/875 + 770/875 = 1220/875
1220/875 = 1 345/875
345/875
Divide each side by 5
69/175
1 + 69/175 = 1 69/175
Answer:
1 69/175